# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ 4.3 Additive Bi-valued Costs
In this part, we consider additive bi-valued cost functions.For any agent $i\in N$, her cost function $c_{i}$ is additive bi-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$.Similar to the results of goods setting, even for two agents, an EF1 or MMS allocation cannot be guaranteed.
@@ 4.3.1 Two Agents
First, we study the additive bi-valued cost functions for two agents and propose the Adapted Chores Envy-Graph Procedures (Algorithm 8) to return an allocation that achieves the tight approximation guarantee for EF1 and has a good approximation for MMS. Using techniques similar to the case of goods, we take the chores adapted envy-graphs procedures as the base algorithm, and then design two auxiliary algorithms (Algorithms 9 and 10) to remove the possible envy cycle by complicated case analysis to ensure the approximation of EF1 and MMS.
Definition 11 (Envy Graph for Chores).
Given an allocation $A^{k}$, in an envy graph, each vertex corresponds to an agent, and each directed edge $(i,j)$ represents that agent $i$ envies agent $j$, where we have $c_{i}(A_{i}^{k})>c_{i}(A_{j}^{k})$.
Input: An instance $(T,N,(c_{1},c_{2}))$ with additive bi-valued cost functions
Output: A complete approximate EF1 and MMS allocation $A$
@@
2Let $A=(\emptyset,\ldots,\emptyset)$ and build an envy graph $G$ for two agents;
@@
4
5      if there is no envy between the two agents then
6             if $\forall k\in N$, $c_{k}(e)=a_{k}$ or $\forall k\in N$, $c_{k}(e)=b_{k}$  then
7                   Arbitrarily choose one agent $i\in N$;
@@
9                   Choosing the agent $i\in N$ who values it at $a_{i}$;
@@
13             Let $i\in N$ be an agent who does not envy the other agent and $j=2-i$;
14            if $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ then
15                   Terminate this algorithm and implement the chores preliminary-breaking-cycle algorithm (Algorithm 9);
16
@@ 18      $A_{i}=A_{i}\cup\{e\}$;
19
return $A$;
Algorithm 8 Adapted Chores Envy-Graph Procedures
Input: An instance $(T,N,(c_{1},c_{2}))$ with additive bi-valued cost functions
Output: A complete approximate EF1 and MMS allocation $A$
@@
2Initialize: $\lambda=1$ and $\mu=0$;
3 when item $e\in T$ arrives do
4
5      if $\lambda=1$ then
6
7            if $|A_{j}|=1$ then
@@ 9                  $A_{j}=A_{j}\cup\{e\}$;
10                  Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
@@ 19                  $\lambda=2$;
21
22      else
23
24            if $\mu=1$  then
25
@@ 26                  $A_{i}=A_{i}\cup\{e\}$;
27            else
28
29                  if $c_{j}(e)=a_{j}$ then
30
31                        $A_{j}=A_{j}\cup\{e\}$;
32                  else
33
34                        $A_{i}=A_{i}\cup\{e\}$;
35                        Terminate this algorithm and execute the chores deep-breaking-cycle algorithm (Algorithm 10);
36                   end if
37
38             end if
39
40       end if
41      Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
42
return $A$;
Algorithm 9 Chores Preliminary-Breaking-Cycle Algorithm
Input: An instance $(T,N,(c_{1},c_{2}))$ with additive bi-valued cost functions
Output: A complete approximate EF1 and MMS allocation $A$
1 Initialize: $\lambda=1$ and $\mu=0$;
2 when item $e\in T$ arrives do
3
4      if $\lambda=1$ then
5
6            if $c_{j}(e)=a_{j}$ and $c_{i}(e)=b_{i}$ then
7                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
8             else if $c_{j}(e)=b_{j}$ and $c_{i}(e)=b_{i}$ then
9
10                  $A_{j}=A_{j}\cup\{e\}$ and $\mu=2$;
11             else if  $c_{j}(e)=a_{j}$ and $c_{i}(e)=a_{i}$  then
12                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=3$;
13             else
14                   $A_{i}=A_{i}\cup\{e\}$ and $\mu=4$;
15             end if
16            $\lambda=2$;
17      else
18             if $\mu=1$ then
19                   $A_{j}=A_{j}\cup\{e\}$;
20             else if  $\mu=2$  then
21                   if $c_{j}(e)=b_{j}$ then
22
23                        $A_{i}=A_{i}\cup\{e\}$;
24                        $\lambda=1$;
@@ 32                        $A_{j}=A_{j}\cup\{e\}$;
35                        $A_{i}=A_{i}\cup\{e\}$;
36                        $\lambda=1$;
37                   end if
38
39             else
40
41                  $A_{j}=A_{j}\cup\{e\}$;
42                  $\lambda=1$;
43             end if
44
45            if $\lambda=2$ then
46                   Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
47
48       end if
49
50
51return $A$;
Algorithm 10 Chores Deep-Breaking-Cycle Algorithm
@@ 46                   Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
For the deterministic allocation of indivisible chores for two agents with additive bi-valued cost functions, the Adapted Chores Envy-Graph Procedures (Algorithm 8) computes a complete allocation that satisfies $2$-EF1 and $\frac{5}{3}$-MMS.
Similar to the goods setting, we first show the following key observations and lemmas before the proof of Theorem 12.
@@
When Algorithm 9 is executed, for the first arriving item $e$, we have $c_{j}(e)=a_{j}$.
@@
When Algorithm 9 is ready to be executed, it means that if the arriving item $e$ is allocated to agent $i$, there will be an envy cycle, i.e, $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ and $c_{i}(A_{i}\cup\{e\})>c_{i}(A_{j})$ For agent $j$, her bundle is EF1. Suppose that $c_{j}(e)=b_{j}$. Then, we have $c_{j}(A_{j})\leq c_{j}(A_{i})+c_{j}(f)\leq c_{j}(A_{i})+b_{j}=c_{j}(A_{i}\cup\{e\})$ for some $f\in A_{j}$, which is a contradiction. Therefore, it must hold that $v_{j}(e)=a_{j}$. ∎
@@
When Algorithm 9 is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
@@
Algorithm 9 computes a complete allocation that is EF1 with or without an envy cycle, or 2-EF1.
@@
When Algorithm 9 is executed, the input allocation $A=(A_{1},A_{2})$ is EF1. According to the description of our algorithm, it is executed in at most two rounds. Then, let us consider different rounds.
$\lambda=1$. When the first item $e_{1}$ arrives, we have $c_{j}(e_{1})=a_{j}$ by Observation 4. When $|A_{j}|=1$, we have $c_{j}(A_{j})=b_{j}$ since she envies agent $i$. Then, this item is allocated to agent $j$. For agent $i$, her bundle is EF since she does not envy agent $j$. For agent $j$, her new bundle is EF1 since $|A_{j}|=1$ and $c_{j}(e_{1})=a_{j}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. When $|A_{j}|>1$, let us consider two subcases about the valuation of $e_{1}$.
  • If $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\mu=1$. For agent $j$, her new bundle is $2$-EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(A_{i})\geq c_{j}(e_{1})=a_{j}$. For agent $i$, her bundle is EF since she does not envy agent $j$. Thus, this allocation is $2$-EF1.
  • If $c_{j}(e_{1})=a_{i}$, agent $i$ picks it and we set $\mu=2$. For agent $i$, her new bundle is EF1 since after removing $e_{1}$ from her bundle, we have $c_{j}(A_{j})\leq c_{j}(A_{i})$. For agent $j$, her bundle is EF1. Thus, this allocation is EF1 with an envy cycle.
$\lambda=2$. When the second item $e_{2}$ arrives, let us consider two cases about $\mu$.
  • If $\mu=1$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $c_{i}(A_{i}\cup\{e_{2}\})\leq c_{i}(A_{i})+b_{i}\leq c_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her bundle is EF1 since we have $v_{j}(A_{j}\setminus\{f\})\leq v_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(e_{2})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • If $\mu=2$, let us consider two subcases about the cost of $e_{2}$. When $c_{j}(e_{2})=a_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i})\leq c_{i}(A_{j})$ and $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{2})=a_{j}\leq c_{j}(e_{1})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. When $c_{j}(e_{2})=b_{j}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i})\leq c_{i}(A_{j})$ and $c_{i}(e_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her bundle is EF1 since she does not pick any item in these two rounds. Thus, this allocation is 2-EF1, and the algorithm is terminated.
@@
When Algorithm 10 is executed, there are two properties about allocation $A$:
@@
  • Let $e_{1}$ and $e_{2}$ denote the two new allocated items in Algorithm 9, we have $c_{i}(A_{i}\setminus\{e\}\leq 2c_{i}(A_{j})$, where $e\in\arg\max_{f\in\{e_{1},e_{2}\}}c_{i}(f)$.
@@
When Algorithm 10 is executed, in Algorithm 9 we have $\lambda=2$ and $\mu=2$. By the proof of Lemma 4, it is easy to check the statements are correct. ∎
@@
Algorithm 10 computes a complete allocation that is EF1 with or without an envy cycle, or 2-EF1.
@@
When Algorithm 10 is executed, by Observation 6, the input allocation $A=(A_{1},A_{2})$ is 2-EF1, where $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$ and $c_{i}(A_{i}\setminus\{e\})\leq 2c_{i}(A_{j})$ for some $e\in A_{i}$. Let $f_{1}$ and $f_{2}$ denote the two online items that are allocated to agent $j$ in Algorithm 9 and we have $\min\{c_{j}(e_{1}),c_{j}(e_{2})\}=a_{j}$.
@@
$\lambda=1$. When the item $e_{1}$ arrives at $\lambda=1$, let us consider four cases of agent’s costs functions.
  • When $c_{j}(e_{1})=a_{j}$ and $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\mu=1$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(f_{1})$. Thus, this allocation is EF1 with an envy cycle.
  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\mu=2$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=b_{j}\leq c_{j}(f_{2})$. Thus, this allocation is EF1 with an envy cycle.
  • When $c_{j}(e_{1})$ and $c_{i}(e_{1})=a_{i}$, agent $j$ picks it and we set $\mu=3$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is 2-EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{i}\leq c_{j}(f_{1})$. Thus, this allocation is EF1 with an envy cycle.
  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=a_{j}$, agent $i$ picks it and we set $\mu=4$. For agent $i$, her new bundle is still 2-EF1 since we have $\frac{c_{i}(A_{i}\setminus\{f\}\cup\{e_{1}\})}{c_{i}(A_{j})}\leq\frac{c_{i}(A_{j})+c_{i}(e_{1})}{c_{i}(A_{j})}\leq 2$ for some $f\in\{f_{1},f_{2}\}$ where the inequality follows from the fact that $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}$ and $|A_{j}|\geq 2$. Thus, this allocation is 2-EF1.
$\lambda=2$. When the item $e_{2}$ arrives at $\lambda=2$, let us consider four cases about $\mu$.
@@
  • When $\mu=1$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{2})\leq c_{i}(e_{1})=b_{i}$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$, $c_{j}(e_{1})=a_{j}$ and $c_{j}(e_{2})\leq c_{j}(f_{2})=b_{j}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • When $\mu=2$, we need to consider two subcases. If $c_{j}(e_{2})=b_{j}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$ and $c_{j}(e_{2})\leq c_{i}(e_{1})=b_{i}$. For agent $j$, her bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{1})=b_{j}=c_{j}(e_{2})$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda=1$ and continue executing this algorithm. If $c_{j}(e_{2})=a_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{2})\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{j}(e_{1})=b_{j}=c_{j}(f_{2})$ and $c_{j}(e_{2})=a_{j}=c_{j}(f_{1})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.

### v2

@@ 4.3 Additive Personalized Bi-valued Costs
In this part, we consider additive personalized bi-valued cost functions. For any agent $i\in N$, her cost function $c_{i}$ is additive personalized bi-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$. Similar to the results of goods setting, even for two agents, an EF1 or MMS allocation cannot be guaranteed.
@@ 4.3.1 Two Agents
First, we study the additive personalized bi-valued cost functions for two agents and propose the Adapted Chores Envy-Graph Procedure (Algorithm 9) to return an allocation that achieves the tight approximation guarantee for EF1 and has a good approximation for MMS. Using techniques similar to the case of goods, we take the Chores Envy-Graphs procedure (Algorithm 10) as the base algorithm. When it may not be further executed (an envy cycle will appear), we skip to the Chores Preliminary-Breaking-Cycle algorithm (Algorithm 11), where we carefully allocate the future items and guarantee the desired approximation. During the execution of the Chores Preliminary-Breaking-Cycle (CPBC) algorithm, in one complicated case ($(\lambda_{1},\mu_{1})=(2,2)$), we introduce the Chores Deep-Breaking-Cycle algorithm (Algorithm 12) to address it. The Chores Deep-Breaking-Cycle (CDBC) algorithm may be consecutively executed in many iterations, but the CPBC algorithm is at most consecutively executed in two iterations. Note that the CPBC and CDBC algorithms are used to remove the possible envy cycle and guarantee that in some further iteration, we can return to the CEGP algorithm.
Input: An instance $(T,N,(c_{1},c_{2}))$ with additive personalized bi-valued cost functions
Output: A complete approximate EF1 and approximate MMS allocation $\bf A$}
@@
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
Initialize $\alpha=1$; // Algorithm Indicator
3
Initialize $(\lambda_{1},\mu_{1})$ and $(\lambda_{2},\mu_{2})$, where $\lambda_{1}=\lambda_{2}=1$ and $\mu_{1}=\mu_{2}=0$; ;
  // $\lambda$: the times of consecutive execution for the CPBC or CDBC algorithm, and $\mu$: the case label in the CPBC or CDBC algorithm.
4
5Build an envy graph $G$ for two agents;
@@
7 if $\alpha=1$ then
8
9    $A,\alpha\leftarrow\mathsf{CEGP}(N,A,\alpha,e)$;
10    if $\alpha=2$ then
       $A,\alpha\leftarrow\mathsf{CPBC}(N,A,\alpha,e,(\lambda_{1},\mu_{1}))$; ;
          // Envy cycle will appear.
11
12      end if
13
14 else if $\alpha=2$ then
15    $A,\alpha,(\lambda_{1},\mu_{1})\leftarrow\mathsf{CPBC}(N,A,\alpha,e,(\lambda_{1},\mu_{1}))$;
@@
17    $A,\alpha,(\lambda_{2},\mu_{2})\leftarrow\mathsf{CDBC}(N,A,\alpha,e,(\lambda_{2},\mu_{2}))$;
@@
19
20  Update the envy graph $G$;
21
return $\bf A$};
Algorithm 9 (ACEGP) Adapted Chores Envy-Graph Procedure
Input: $N,{\bf A},\alpha,e$
Output: A complete allocation $\bf A$} and $\alpha$
1
2if  there is no envy between the two agents  then
3 if $\exists i\in N$, $c_{i}(e)=a_{k}$  then
4      Choosing the agent $i$ who values it at $a_{i}$ (breaking ties arbitrarily);
5 else
6      Arbitrarily choose one agent $i$;
7   end if
8
9else
10   Let $i\in N$ be an agent who does not envy the other agent and j=2ij=2-i;
11 if $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ then
12    $\alpha=2$;
13    return $\bf A$} and $\alpha$;
14   end if
15
16 end if
[…]
return $\bf A$} and $\alpha$;
Algorithm 10 (CEGP) Chores Envy-Graph Procedure
Input: $N,{\bf A},\alpha,e,(\lambda_{1},\mu_{1})$
Output: A complete allocation $\bf A$}, $\alpha$ and $(\lambda_{1},\mu_{1})$
1
2if $\lambda_{1}=1$ then
3
4 if $|A_{j}|=1$ then
5
6    $A_{j}=A_{j}\cup\{e\}$;
7    $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$;
8 else
9
10    if $c_{i}(e)=b_{i}$ then
@@ 13       $\lambda_{1}=2$ and $\mu_{2}=1$;
15       $A_{i}=A_{i}\cup\{e\}$;
16       $\lambda_{1}=2$ and $\mu_{2}=2$;
@@ 25    $A_{i}=A_{i}\cup\{e\}$;
26 else
27
28    if $c_{j}(e)=a_{j}$ then
29
30       $A_{j}=A_{j}\cup\{e\}$;
31    else
32
33       $A_{i}=A_{i}\cup\{e\}$;
34       $\alpha=3$, $\lambda_{1}=1$, and $\mu_{2}=0$;
35      end if
36
37   end if
38
39 $\alpha=1$, $\lambda_{1}=1$, and $\mu_{2}=0$;
40 end if
41
return $\bf A$}, $\alpha$ and $(\lambda_{1},\mu_{1})$;
Algorithm 11 (CPBC) Chores Preliminary-Breaking-Cycle
Input: $N,{\bf A},\alpha,e,(\lambda_{2},\mu_{2})$
Output: A complete allocation $\bf A$}, $\alpha$, and $(\lambda_{2},\mu_{2})$
@@ 39 $\alpha=1$, $\lambda_{1}=1$, and $\mu_{2}=0$;
2if $\lambda_{2}\mod 2=1$ then
3
4 if $c_{j}(e)=a_{j}$ then
5    $A_{j}=A_{j}\cup\{e\}$;
6    $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$;
7 else if $c_{j}(e)=b_{j}$ and $c_{i}(e)=b_{i}$ then
@@ 9    $A_{j}=A_{j}\cup\{e\}$;
10    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=1$;
@@ 13    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=2$;
17else if $\lambda_{2}\mod 2=0$ then
18 if  $\mu_{2}=1$  then
19    if $c_{j}(e)=b_{j}$ then
20
@@ 21       $A_{i}=A_{i}\cup\{e\}$;
22       $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
@@ 31    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
35return $\bf A$}, $\alpha$, and $(\lambda_{2},\mu_{2})$;
Algorithm 12 (CDBC) Chores Deep-Breaking-Cycle
Definition 11 (Envy Graph for Chores).
Given an allocation ${\bf A}^{k}$, the corresponding envy graph is defined as $G=(V,E)$, where the vertex set $V$ corresponds to agent set $N$, and a directed edge $(i,j)\in E$ iff agent $i$ envies agent $j$, i.e., $c_{i}(A_{i}^{k})>c_{i}(A_{j}^{k})$. Additionally, the directed cycle in the envy graph is called an envy cycle.
@@
For the deterministic allocation of indivisible chores for two agents with additive personalized bi-valued cost functions, the Adapted Chores Envy-Graph Procedures (Algorithm 9) computes a complete allocation that satisfies $2$-EF1 and $\frac{5}{3}$-MMS.
Before we show the above Theorem, we first list some lemmas used in the proof.
Lemma 8.
The CEGP algorithm (Algorithm 10) computes an EF1 allocation without an envy cycle.
Proof.
Note that there are three kinds of input allocation. The first one is the initial empty allocation, which is EF1 without an envy cycle. The second one is the output allocation of the CPBC algorithm. The last one is the output allocation of the CDBC algorithm. It suffices to show that the last two kinds of input allocation are EF1 without an envy cycle. That is because if the input allocation of the CEGP algorithm is EF1 without an envy cycle, by the description of the algorithm and induction, it is easy to check that the property of EF1 without an envy cycle is always maintained. By Lemmas 9 and 10, we can derive that when the CPBC or CDBC algorithm outputs $\alpha=1$, implying that the CEGP algorithm is ready to be executed, the output allocation of the CPBC or CDBC algorithm is EF1 without an envy cycle. Therefore, our claim holds. ∎
@@
When the CPBC algorithm (Algorithm 11) is executed, for the first arriving item $e$, we have $c_{j}(e)=a_{j}$.
@@
When the CPBC algorithm is ready to be executed, it means that if the arriving item $e$ is allocated to agent $i$, there will be an envy cycle, i.e, $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ and $c_{i}(A_{i}\cup\{e\})>c_{i}(A_{j})$ For agent $j$, her bundle is EF1. Suppose that $c_{j}(e)=b_{j}$. Then, we have $c_{j}(A_{j})\leq c_{j}(A_{i})+c_{j}(f)\leq c_{j}(A_{i})+b_{j}=c_{j}(A_{i}\cup\{e\})$ for some $f\in A_{j}$, which is a contradiction. Therefore, it must hold that $v_{j}(e)=a_{j}$. ∎
@@
When the CPBC algorithm (Algorithm 11) is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
@@
The CPBC algorithm (Algorithm 11) computes a complete allocation satisfying:
  • 2-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
@@
The input allocation $\textbf{A}=(A_{1},A_{2})$ of the CPBC algorithm is EF1. According to the description of our algorithm, it is executed in at most two rounds. Then, let us consider different rounds.
$\lambda_{1}=1$. When the first item $e_{1}$ arrives, we have $c_{j}(e_{1})=a_{j}$ by Observation 4. When $|A_{j}|=1$, we have $c_{j}(A_{j})=b_{j}$ since she envies agent $i$. Then, this item is allocated to agent $j$. For agent $i$, her bundle is EF since she does not envy agent $j$. For agent $j$, her new bundle is EF1 since $|A_{j}|=1$ and $c_{j}(e_{1})=a_{j}$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that the CEGP algorithm is ready to be implemented. When $|A_{j}|>1$, let us consider two subcases about the cost of $e_{1}$.
  • If $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\lambda_{2}=2$ and $\mu_{1}=1$. For agent $j$, her new bundle is $2$-EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(A_{i})\geq c_{j}(e_{1})=a_{j}$. For agent $i$, her bundle is EF since she does not envy agent $j$. Thus, this allocation is $2$-EF1.
  • If $c_{j}(e_{1})=a_{i}$, agent $i$ picks it and we set $\lambda_{2}=2$ and $\mu_{2}=2$. For agent $i$, her new bundle is EF1 since after removing $e_{1}$ from her bundle, we have $c_{j}(A_{j})\leq c_{j}(A_{i})$. For agent $j$, her bundle is EF1. Thus, this allocation is EF1 with an envy cycle.
In the above cases, the CPBC algorithm continues to be executed.
$\lambda_{1}=2$. When the second item $e_{2}$ arrives, let us consider two cases about $\mu_{1}$.
  • If $\mu_{1}=1$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $c_{i}(A_{i}\cup\{e_{2}\})\leq c_{i}(A_{i})+b_{i}\leq c_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her bundle is EF1 since we have $v_{j}(A_{j}\setminus\{f\})\leq v_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(e_{2})$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that the CEGP algorithm is ready to be executed.
  • If $\mu_{1}=2$, let us consider two subcases about the cost of $e_{2}$. When $c_{j}(e_{2})=a_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i})\leq c_{i}(A_{j})$ and $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{2})=a_{j}\leq c_{j}(e_{1})$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that the CEGP algorithm is ready to be executed. When $c_{j}(e_{2})=b_{j}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i})\leq c_{i}(A_{j})$ and $c_{i}(e_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her bundle is EF1 since she does not pick any item in these two rounds. Thus, this allocation is 2-EF1, and we set $\alpha=3$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that we need to implement the CDBC algorithm.
@@
When the CDBC algorithm (Algorithm 12) is executed, there are two properties about allocation A:
@@
  • Let $e_{1}$ and $e_{2}$ denote the two new allocated items in the CPBC algorithm, we have $c_{i}(A_{i}\setminus\{e\}\leq 2c_{i}(A_{j})$, where $e\in\arg\max_{f\in\{e_{1},e_{2}\}}c_{i}(f)$.
@@
When the CDBC algorithm is executed, the CPBC algorithm is terminated for $\lambda_{1}=\mu_{1}=2$. Then, by the proof of Lemma 9, it is easy to check that statements are correct. ∎
@@
The CDBC algorithm (Algorithm 12) computes a complete allocation satisfying:
  • 2-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
@@
When the CDBC algorithm is executed, by Observation 6, the input allocation $A=(A_{1},A_{2})$ is 2-EF1, where $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$ and $c_{i}(A_{i}\setminus\{e\})\leq 2c_{i}(A_{j})$ for some $e\in A_{i}$. Let $f_{1}$ and $f_{2}$ denote the two items that are allocated to agent $j$ in the CPBC algorithm and we have $\min\{c_{j}(e_{1}),c_{j}(e_{2})\}=a_{j}$.
@@
$\lambda_{2}=1$. When the item $e_{1}$ arrives at $\lambda_{2}=1$, let us consider three cases of cost functions.
@@
  • When $c_{j}(e_{1})=a_{j}$, agent $j$ picks it and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF since we have $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$. Thus, this allocation is EF1 without an envy cycle, implying that the CEGP algorithm is ready to be executed.